A reaction between dinitrogen tetrahydride and dinitrogen tetroxide has been used to launch rockets into space. The reaction produced water vapor and nitrogen gas. How many molecules are released into the atmosphere of 450kg of the oxide are consumed?


1
Expert's answer
2021-01-06T05:58:06-0500

Solution:


Chemical reaction:

"2N_2H_4+N_2O_4\rightarrow 3N_2+4H_2O"


Total molecules released into the air:


"450,000\cdot\;\dfrac{1\;mole}{92}\cdot 6.02*10^{23}" = "2.94*10^{26}\;molecules"


Answer:

"2.94*10^{26}\;molecules"